OpenStudy (anonymous):

If Ken wanted to create a function that modeled a base of 11 and what exponents were needed to reach specific values, how would he set up his function?

OpenStudy (anonymous):

\[a. f(x)=11^x\] \[b. f(x)=x^11\] \[c. f(x)=\log _{11}^{x}\] \[d. f(x)=\log _{x}11\]

OpenStudy (anonymous):

I think the answer is D....

OpenStudy (anonymous):

@bhowl77

OpenStudy (anonymous):

C has a log base 11 so I'd do that...but then I'm not sure I understand the question.

OpenStudy (anonymous):

nope, im thinking A now.

OpenStudy (anonymous):

yeah, I think its A. That's a base 11 raised to an exponent of x, so that fits the question pretty well.

OpenStudy (anonymous):

:/ I don't get logarithms at all This question seems easier. Express \[125=5^x\] as a logarithmic equation

OpenStudy (anonymous):

\[a. \log_5^x=25\]

OpenStudy (anonymous):

logbase5(125) is 3... i think

OpenStudy (anonymous):

so log5(x)=3 Its been forever since I did this stuff. so I'm a bit shaky on that.

OpenStudy (anonymous):

like this \[\log_5125=x\]

OpenStudy (anonymous):

yeah, and log5(125) is 3. so x=3

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

np